6 cm Session

Session: Nov 2003 (session 3)
Wavelength: 6 cm


To give feedback on this session, click here

6 cm Session Feedback

Jodrell_Bank (Alastair Gunn at Fri Nov 7 02:45:04 2003) Failure :
Whilst setting up for L-band we found a fault in our final LO synthesiser which has probably compromised most of our observations at 5GHz this session. The synthesiser wasn't producing the frequency it was set to, but may still have been locked. The difference is around 200KHz, which may be too large to compensate for at correlation. We are using an alternative synthesiser for L-band observations.

JIVE (Bob Campbell at Mon Jan 19 11:33:14 2004) Failure :
Further investigation of Jb data showed that the frequency of phase-cal tones was unstable on the order of a couple Hz on the timescale of seconds. Use of an LO-offset determined from a time-averaged value of the offset of the phase-cal tones from their expected frequencies resulted in only intermittent fringes on baselines to Jb. Both these findings are consistent with the frequency not being sufficiently locked to correlate reliably. Similar behavior was also seen in Socorro. Therefore we dropped Jb for the 6cm experiments from this session.